# A Distributed ADMM-based Deep Learning Approach for Thermal Control in Multi-Zone Buildings

### v1

@@ Abstract
The surge in electricity use, coupled with the dependency on intermittent renewable energy sources, poses significant hurdles to effectively managing power grids, particularly during times of peak demand. Demand Response programs and energy conservation measures are essential to operate energy grids while ensuring a responsible use of our resources This research combines distributed optimization using ADMM with Deep Learning models to plan indoor temperature setpoints effectively. A two-layer hierarchical structure is used, with a central building coordinator at the upper layer and local controllers at the thermal zone layer. The coordinator must limit the building’s maximum power by translating the building’s total power to local power targets for each zone. Local controllers can modify the temperature setpoints to meet the local power targets. The resulting control algorithm, called Distributed Planning Networks, is designed to be both adaptable and scalable to many types of buildings, tackling two of the main challenges in the development of such systems. The proposed approach is tested on an 18-zone building modeled in EnergyPlus. The algorithm successfully manages Demand Response peak events.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Buildings account for approximately $40\%$ of the energy consumption worldwide. More than half of the building’s energy is used to maintain the occupant’s comfort, mainly through Heating, Ventilation and Air Conditioning (HVAC) systems [1]. Buildings are powerful pillars for integrating energy efficiency and Demand Response (DR) measures to reduce electricity consumption, especially during unstable power supply and peak periods. Many paths are investigated to benefit grid management by leveraging the building’s power consumption. Local Renewable Energy Sources, Energy Storage Systems, and Electric Vehicles can now be embedded in the Building Energy Management System (BEMS) to form a building-integrated microgrid [2, 3]. Efforts are also made to improve insulation materials and directly reduce the energy needs for heating and cooling purposes [4]. In addition, many grid operators are now proposing DR programs to encourage consumers to reduce their energy consumption following an incentive-based or dynamic pricing program. In both cases, they propose rewards to the end-users in exchange for reductions in their power or energy consumption, especially during periods of high solicitation of the grid. Thus, smart control of the building loads may result in substantial financial advantages for the consumers while providing support for an efficient grid operation [5, 6, 7]. Heating and cooling demands can significantly impact the total building’s power consumption profile. The thermal inertia of the building offers enhanced operational flexibility to reduce/shift a part of the load using pre-cooling or pre-heating strategies. The use of smart temperature controllers, accounting for occupancy and weather variations, results in a direct reduction of energy usage. However, it has a direct impact on the user’s comfort that must be preserved [5]. Many works have focused on the creation of advanced BEMS for monitoring and controlling a building’s energy requirements, including smart temperature controllers. Two of the main approaches used to develop BEMS are Model Predictive Control (MPC) and Reinforcement Learning (RL). For instance, [8, 9, 10] apply MPC to manage energy consumption and comfort in commercial and residential buildings. Works such as [11, 12, 13, 14, 15] use RL to create efficient and adaptive controllers for BEMS.
@@
Optimization problems solved using MPC or RL are often formulated as centralized problems. In centralized problems, a single entity computes the solution to the problem and takes all the decisions. Centralized problem formulations for BEMS are further described in [16]. Although simple conceptually, this approach is not scalable as the system size grows [17]. Furthermore, the necessary communications between local actuators and the central controller may raise privacy concerns and failure risks. Thus, distributed architectures are increasingly considered. Distributed architectures are of two types: the hierarchical approach and the fully distributed approach (peer-to-peer). In both approaches, the complex global optimization problem is decomposed into subproblems. In hierarchical structures, the problem is solved through a coordinator who takes global decisions and communicates with Local Controllers (LCs). In fully distributed architectures, LCs directly communicate with each other to coordinate their actions. The choice of architecture is problem dependent, as each architecture has pros and cons and highly depends on the communication infrastructure [18, 11].
@@
On the theoretical side, ADMM is proven to converge when the underlying problem is convex [24], which is not the case in BEMS applications. To rely on the convergence guarantees of convex problems, many works linearize the dynamics of their system [19, 20, 25, 26, 27, 28]. Others rely on the empirical robustness of ADMM when applied to non-convex problems [22]. Few theoretical results exist on the convergence of ADMM in non-convex settings [29, 30]. The use of a non-convex ADMM allows to consider models more sophisticated than linear ones. In particular, models relying on Neural Networks proved to be extremely accurate to forecast the energy consumption in buildings [31, 23]. Recurrent State Space Models (RSSM) also proved to be extremely accurate to model stochastic environments and are yet to be applied to building energy forecasting. Such models, along with planning algorithms provide state-of-the-art results in many benchmark control tasks [32, 33].
@@
In line with the International Energy Agency’s (IEA) recommendations [34], the present study focuses on the critical role of widening temperature setpoints as a way for significant energy savings and effective implementation of DR programs. While optimizing the HVAC system’s operation to efficiently track a setpoint is important, it is crucial to note that for a given temperature setpoint, there exists an inherent physical constraint on the extent of energy reduction achievable solely through the optimization of HVAC system operations. Further energy reduction may only be achieved by altering the setpoint itself. This study thus uses temperature setpoints as actuators for power consumption regulation. Within the proposed framework, users select a preferred temperature setpoint and define an associated comfort range for each zone within a building. The BEMS dynamically adjusts the setpoints within the comfort intervals to distribute energy savings across the building.
@@
Our approach combines the recent advancement in non-convex ADMM algorithms, with state-of-the-art control methods based on Deep Learning models. First, we formulate the temperature control problem as a centralized problem with a constraint on the building’s total power consumption. Then, we convert this problem into a non-convex consensus problem that we decompose using the ADMM algorithm. We provide convergence proof of the algorithm for this non-convex setting, which is key to accurately account for the non-linear dynamics of buildings. From this formulation, we derive a surrogate problem that can be efficiently solved in practice. We test both RSSM and SSM to model the building’s zone. The models are used to plan for the temperature setpoints that best match the comfort requirements of zones while enforcing a limit on the building’s power usage. The two models are compared in terms of prediction accuracy to benchmark RSSM in a building environment, and in terms of control quality to assess the performance of our distributed control framework. The main contributions of this study are summarized as follows:
@@
    This work proposes an efficient algorithm to adjust a building’s indoor temperature based on available energy resources. The approach contributes to energy conservation while minimally affecting occupant comfort, making it a valuable asset for DR.
@@
    We combine non-convex ADMM and advanced, fully data-driven, Deep Learning models to provide a distributed control for HVAC systems. To the best of our knowledge, this is the first work combining non-linear Deep Learning models and ADMM.
@@
    While the global constraint is on the power consumption of the building, we propose to use temperature setpoints of the zones as actuators. Therefore, no assumption is made on the type of HVAC system, which makes our approach easily adaptable from one building to another.
  4. 4.
    The building is modeled using a fully data-driven approach, where no prior knowledge about the building’s blueprint is required. The RSSM allows one to plan in a stochastic environment and handle uncertainties in a robust MPC framework. The SSM can be used in computationally efficient shooting controllers.
  5. 5.
    We propose a convergence analysis for our non-convex ADMM algorithm. A case study shows the good convergence properties of our algorithm in practice.
II Problem formulation
In this study, a building composed of $N$ thermal zones is considered. Each zone has a temperature control system, i.e. a device like a PID or ON/OFF controller able to track a local temperature setpoint. Specifically, we consider a discretized setting, where occupants choose temperature setpoints $T^{sp}_{t,i}$ for each zone, that are optimal for their comfort. These setpoints may vary in time but are known beforehand. At time $t$ for the zone $i$, $u_{t,i}$ represents the power consumption of a Heating or Cooling (HC) system.
The power consumption is closely related to the temperature setpoints [35]. One can act on the setpoints to control the building’s HC power consumption below a maximum value. To this purpose, the global power constraint must be distributed to local temperature controllers acting on each zone, and translated to temperature setpoints.
The proposed control system modifies the occupants’ setpoints $T^{sp}_{t,i}$ by $\delta_{t,i}$, so that the setpoint tracked by the HC system is $T^{sp}_{t,i}+\delta_{t,i}$ at time $t$ is zone $i$. Let
@@ II Problem formulation
The optimal setpoint changes $\boldsymbol{\Delta}$ are solutions to the following optimization problem:
{mini!}
[…]
@@
  • $s_{t,i}$ is the state of zone $i$. The state includes, in particular, $T^{sp}_{t,i}$, the air temperature, and the HC power consumption. The state is partially observable as it may also include the temperature of the walls or the thermal capacity of the zone.
@@
    The coefficient $A_{1,1}$ selects the power consumption of zone $1$ at the first time step.
  • $\textbf{P}^{max}=[P^{max}_{1},\ldots,P^{max}_{H}]$ is a vector containing the maximum power for each timestep. The vector inequality in eq. (II) is component-wise.
  • $M_{i}$ is the maximum setpoint change defined by the occupants in zone $i$.
To solve this problem efficiently, we reformulate it as a sharing problem [29] and apply ADMM.
First, the inequality constraint (II) is relaxed using a new parameter $\textbf{P}^{tot}$ such that $\textbf{P}^{tot}\leq\textbf{P}^{max}$. This ensures the problem is feasible and removes a coupling constraint. Second, the true dynamics $f_{i}$ is approximated by a function $q_{i}$ that does not take the neighboring states as input. This assumption implies that one can reasonably predict a zone’s state without having to predict the states of neighboring zones. Note that the initial state of a zone $i$, $s_{0,i}$ may still be augmented to include information about the neighbors at $t=0$. In many buildings such as hotels, offices, or residential buildings, the temperature ranges in the zones are similar, thus limiting the heat exchanges. We will assess this assumption experimentally in section V to show that it is reasonable if the zones are well insulated. Such an assumption is also discussed in [22, 36]. The optimization problem becomes:
{mini!}
[…]
The HC power consumption is now driven towards the parameters $\textbf{P}^{tot}=[P_{1}^{tot},\ldots,P_{H}^{tot}]$. The choice of $\textbf{P}^{tot}$ is important and will be further discussed in section V.
Let us define a function $\phi_{i}$ that maps the state of a zone $i$ to its power usage:
$u_{t,i}=\phi_{i}(s_{t,i}).$
Given the setpoints and the first state, future states are computed by unrolling the dynamics:
$s_{1,i}$ $=q_{i}(s_{0,i},\delta_{0,i})$
$s_{2,i}$ $=q_{i}(s_{1,i},\delta_{1,i})$
$=q_{i}(q_{i}(s_{0,i},\delta_{0,i}),\delta_{1,i})$
$\ldots$
$s_{H,i}$ $=q_{i}(s_{H-1,i},\delta_{H,i})$
$=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{H,i})$
To ease the notations, we define the following functions, for all $i$ and $t$:
$Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{t,i})$ (1)
@@
$\ell(\boldsymbol{\Delta})=\sum_{t=1}^{H}\left(\sum_{i=1}^{N}\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})-\textbf{P}^{tot}_{t}\right)^{2}$ (3)
The index to the function $Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})$ indicates that it represents the state at time $t$ of zone $i$ computed from $s_{0,i}$. Only the first $t-1$ elements of $\boldsymbol{\Delta}_{i}$ are needed to compute $s_{t,i}$, but the entire vector is passed for notational simplicity. The power usage becomes:
$u_{t,i}=\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i}))$ (4)
[…]
This problem is a sharing problem as described in [29]. Each zone has a local objective represented by $g$ and a global objective represented by $\ell$. The total power available is a common resource that must be shared across all zones to satisfy the local comfort requirements.
[…]

### v2

@@ Abstract
The increasing electricity use and reliance on intermittent renewable energy sources challenge power grid management during peak demand, making Demand Response programs and energy conservation measures essential. This research combines distributed optimization using ADMM with deep learning models to plan indoor temperature setpoints effectively. A two-layer hierarchical structure is used, with a central building coordinator at the upper layer and local controllers at the thermal zone layer. The coordinator must limit the building’s maximum power by translating the building’s total power to local power targets for each zone. Local controllers can modify the temperature setpoints to meet the local power targets. While most algorithms are either centralized or require prior knowledge about the building’s structure, our approach is distributed and fully data-driven. The proposed algorithm, called Distributed Planning Networks, is designed to be both adaptable and scalable to many types of buildings, tackling two of the main challenges in the development of such systems. The proposed approach is tested on an 18-zone building modeled in EnergyPlus. The algorithm successfully manages Demand Response peak events.
@@ I Introduction
Buildings account for approximately $40\%$ of the energy consumption worldwide. More than half of the building’s energy is used to maintain the occupant’s comfort, mainly through Heating, Ventilation and Air Conditioning (HVAC) systems [1]. Buildings are crucial for integrating energy efficiency and Demand Response (DR) measures to reduce electricity consumption, especially during unstable power supply and peak periods. Many paths are investigated to benefit grid management by leveraging the building’s power consumption. Local renewable energy sources, energy storage systems, and electric vehicles can now be embedded in the Building Energy Management System (BEMS) to form a building-integrated microgrid [2, 3]. Efforts are also made to improve insulation materials and directly reduce the energy needs for heating and cooling purposes [4]. In addition, many grid operators are now proposing DR programs to encourage consumers to reduce their energy consumption following an incentive-based or dynamic pricing program. In both cases, they propose rewards to the end-users in exchange for reductions in their power or energy consumption, especially during periods of high solicitation of the grid. Thus, smart control of the building loads may result in substantial financial advantages for the consumers while providing support for an efficient grid operation [5, 6, 7]. Heating and cooling demands can significantly impact the total building’s power consumption profile. The thermal inertia of the building offers enhanced operational flexibility to reduce/shift a part of the load using pre-cooling or pre-heating strategies. The use of smart temperature controllers, accounting for occupancy and weather variations, results in a direct reduction of energy usage. However, it has a direct impact on the user’s comfort that must be preserved [5]. Many works have focused on the creation of advanced BEMS for monitoring and controlling a building’s energy requirements, including smart temperature controllers. Two of the main approaches used to develop BEMS are Model Predictive Control (MPC) and Reinforcement Learning (RL). For instance, [8, 9, 10] apply MPC to manage energy consumption and comfort in commercial and residential buildings. Works such as [11, 12, 13, 14, 15] use RL to create efficient and adaptive controllers for BEMS.
@@
Optimization problems solved using MPC or RL are often formulated as centralized problems. In centralized problems, a single entity computes the solution to the problem and makes all the decisions. Centralized problem formulations for BEMS are further described in [16]. Although simple conceptually, this approach is not scalable as the system size grows [17]. Furthermore, the necessary communications between local actuators and the central controller may raise privacy concerns and failure risks. Thus, distributed architectures are increasingly considered. Distributed architectures are of two types: the hierarchical approach and the fully distributed approach (peer-to-peer). In both approaches, the complex global optimization problem is decomposed into subproblems. In hierarchical structures, the problem is solved through a coordinator who takes global decisions and communicates with Local Controllers (LCs). In fully distributed architectures, LCs directly communicate with each other to coordinate their actions. The choice of architecture is problem-dependent, as each architecture has pros and cons and highly depends on the communication infrastructure [18, 11].
@@
On the theoretical side, ADMM is proven to converge when the underlying problem is convex [24], which is not the case in BEMS applications. To rely on the convergence guarantees of convex problems, many works linearize the dynamics of their system [19, 20, 25, 26, 27, 28]. Others rely on the empirical robustness of ADMM when applied to non-convex problems [22]. Few theoretical results exist on the convergence of ADMM in non-convex settings [29, 30]. The use of a non-convex ADMM allows to consider models more sophisticated than linear ones. In particular, models relying on neural networks proved to be extremely accurate in forecasting the energy consumption in buildings [31, 23]. Recurrent State Space Models (RSSM) also proved to be extremely accurate in modeling stochastic environments and are yet to be applied to building energy forecasting. Such models, along with planning algorithms provide state-of-the-art results in many benchmark control tasks [32, 33].
@@
In line with the International Energy Agency’s (IEA) recommendations [34], the present study focuses on the critical role of widening temperature setpoints as a way for significant energy savings and effective implementation of DR programs. While optimizing the HVAC system’s operation to efficiently track a setpoint is important, it is crucial to note that for a given temperature setpoint, there exists an inherent physical constraint on the extent of energy reduction achievable solely through the optimization of HVAC system operations. Further energy reduction may only be achieved by modifying the setpoint itself. This study thus uses temperature setpoints as actuators for power consumption regulation. Within the proposed framework, users select a preferred temperature setpoint and define an associated comfort range for each zone within a building. The BEMS dynamically adjusts the setpoints within the comfort intervals to distribute energy savings across the building.
@@
Our approach combines the recent advancement in non-convex ADMM algorithms, with state-of-the-art control methods based on deep learning models. First, we formulate the temperature control problem as a centralized problem with a constraint on the building’s total power consumption. Then, we convert this problem into a non-convex sharing problem that we decompose using the ADMM algorithm. We provide convergence proof of the algorithm for this non-convex setting, which is key to accurately account for the non-linear dynamics of buildings. From this formulation, we derive a surrogate problem that can be efficiently solved in practice. We test both State Space Models (SSM) and RSSM to model the building’s zone. The models are used to plan for the temperature setpoints that best match the comfort requirements of zones while enforcing a limit on the building’s power usage. The two models are compared in terms of prediction accuracy to benchmark RSSM in a building environment, and in terms of control quality to assess the performance of our distributed control framework. The main contributions of this study are summarized as follows:
@@
    We combine non-convex ADMM and fully data-driven deep-learning models to develop a distributed control algorithm for HVAC systems. A convergence analysis of the ADMM algorithm in this non-convex setting is presented. To the best of our knowledge, this is the first work combining non-linear deep-learning models and ADMM.
@@
    We propose a hard constraint of the maximum HVAC power consumption to prioritize energy savings during peak periods, while the use of comfort intervals instead of fixed temperature setpoints guarantees a minimum level of comfort for the users.
@@
    We derive a deterministic and a stochastic version of our algorithm. The performances of both approaches are evaluated on a multi-zone building during DR events. The results demonstrate the control system’s capability to lower power consumption during peak periods. The algorithm consistently converges to a solution with each iteration. Additionally, the distributed architecture of the control system improves its scalability as the building size increases.
In section II, we introduce the control problem and derive a distributed algorithm to solve it. In section III, we present the deep learning models used to model the thermal zones and the implementation of the proposed distributed control algorithm. In section IV, we present experimental results on the prediction accuracy of the models and the control performances of the algorithm. Finally, we discuss the results and some limitations paving the way for future works in section V.
II Description of the Control Problem
II-A Problem Formulation
Consider a building composed of $N$ thermal zones equipped with independent HVAC systems. In each zone, occupants may choose a temperature setpoint $T^{sp}_{t,i}$ that is optimal for their comfort, where $t$ is a time index and $i$ a zone index. These setpoints may vary in time but are known beforehand. To track their setpoint, each HVAC system consumes a given amount of power $u_{t,i}$. The HVAC power consumption is directly impacted by the temperature setpoints [35] and we propose to use the setpoints as actuators to control the HVAC consumption. Specifically, we aim at changing the occupants’ preferred setpoints $T^{sp}_{t,i}$ by $\delta_{t,i}$, so that the HVAC system tracks the setpoint $T^{sp}_{t,i}+\delta_{t,i}$. Let
@@ II-A Problem Formulation
The purpose of the control system is to apply a constraint on the entire building’s HVAC maximum power consumption. We define the optimal setpoint changes $\boldsymbol{\Delta}$ to satisfy this constraint as solutions to the following optimization problem:
$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\|\boldsymbol{\Delta}\|^{2}_{2}$ (1)
subject to $\boldsymbol{s}_{t+1,i}=f_{i}\left(\boldsymbol{s}_{t,i},\delta_{t,i},\boldsymbol{s}_{y,j\in\mathcal{N}_{i}},\delta_{t,j\in\mathcal{N}_{i}}\right),\forall t,i$ (2)
$m_{i}\leq\delta_{t,i}\leq M_{i},\forall t,i$ (3)
$A\textbf{u}\leq\textbf{P}^{max}$ (4)
@@
  • $\mathbb{s}_{t,i}$ is the state of zone $i$. The state includes, in particular, $T^{sp}_{t,i}$, the air temperature, and the HVAC power consumption. Note that the state is partially observable as it may also include the temperature of the walls or the thermal capacity of the zone.
@@
  • $\textbf{P}^{max}=[P^{max}_{1},\ldots,P^{max}_{H}]$ is a vector containing the maximum power for each timestep. The vector inequality in Eq. (4) is component-wise.
  • $m_{i}$ and $M_{i}$ are the minimum and maximum setpoint changes defined by the occupants in zone $i$.
This formulation prioritizes power savings over temperature comfort as long as the temperature setpoint stays in the comfort interval $\left[T^{sp}_{t,i}-m_{i};T^{sp}_{t,i}+M_{i}\right]$ defined by the user.
To solve Problem (1) efficiently, we reformulate it as a sharing problem [29] and apply ADMM to solve it. First, the inequality constraint (4) is relaxed using a new parameter $\textbf{P}^{tot}$ such that $\textbf{P}^{tot}\leq\textbf{P}^{max}$. This ensures the problem is feasible and removes a coupling constraint. Second, the true dynamics $f_{i}$ is approximated by a function $q_{i}$ that does not take the neighboring states as input. This assumption implies that one can reasonably predict a zone’s state without having to predict the states of neighboring zones. Note that the initial state of a zone $i$, $s_{0,i}$ may still be augmented to include information about the neighbors at $t=0$. In many buildings such as hotels, offices, or residential buildings, the temperature ranges in the zones are similar, thus limiting the heat exchanges. We will assess this assumption experimentally in section IV to show that it is reasonable if the zones are well insulated. Such an assumption is also discussed in [22, 36]. The optimization problem becomes:
$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\|\boldsymbol{\Delta}\|^{2}_{2}+||A\textbf{u}-\textbf{P}^{tot}||^{2}_{2}$ (5)
subject to $\boldsymbol{s}_{t+1,i}=q_{i}\left(\boldsymbol{s}_{t,i},\delta_{t,i}\right),\quad\forall t,i$ (6)
$m_{i}\leq\delta_{t,i}\leq M_{i},\forall t,i$ (7)
Note that we do not use weights to balance the two terms of the objective. In practice, the temperature and power values should be normalized for the two terms to have the same order of magnitude. The HVAC power consumption is now driven towards the parameters $\textbf{P}^{tot}=[P_{1}^{tot},\ldots,P_{H}^{tot}]$. The choice of $\textbf{P}^{tot}$ is important and will be further discussed in section III-B3.
The powers u are components of the state s and the constraint Eq. (6) can be made implicit by unrolling the dynamics $q_{i}$ from the initial state using the setpoint schedules. To this purpose, let $\phi_{i}$ map the state of a zone $i$ to its power usage $u_{t,i}=\phi_{i}(\mathbb{s}_{t,i})$, and consider the following functions for all $i$ and $t$:
$Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})=q_{i}(q_{i}(\ldots q_{i}(\mathbb{s}_{0,i},\delta_{0,i})),\delta_{t,i})$ (8)
@@
$\ell(\boldsymbol{\Delta})=\sum_{t=1}^{H}\left(\sum_{i=1}^{N}\phi_{i}(Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})-\textbf{P}^{tot}_{t}\right)^{2}$ (10)
The index to the function $Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})$ indicates that it represents the state at time $t$ of zone $i$ computed from $\mathbb{s}_{0,i}$. Only the first $t$ elements of $\boldsymbol{\Delta}_{i}$ are needed to compute $\mathbb{s}_{t,i}$, but the entire vector is passed for notational simplicity. The power usage of a zone is thus
$u_{t,i}=\phi_{i}(Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})).$ (11)
Using these notations, Problem (5) may be reformulated as a sharing problem
$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\sum_{i=1}^{N}g(\boldsymbol{\Delta}_{i})+\ell(\boldsymbol{\Delta})$ (12)
subject to $m_{i}\leq\delta_{t,i}\leq M_{i}\quad\forall t,i.$ (13)
Each zone has a local objective represented by $g$ and a global objective represented by $\ell$. The total power available is a common resource that must be shared across all zones to satisfy the local comfort requirements.
II-B Distributed Sharing Algorithm
In this section, we derive a distributed algorithm based on ADMM to solve the Problem (12). Note that no particular assumption is made on the dynamics $q_{i}$ of the zones and the problem is thus non-convex. First, we introduce duplicated variables $\bar{\boldsymbol{\Delta}}_{1},\ldots,\bar{\boldsymbol{\Delta}}_{N}$ and form an equivalent problem with $N$ vector linear equality constraints:
$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\sum_{i=1}^{N}g(\boldsymbol{\Delta}_{i})+\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{\Delta}}_{i}\right)$ (14)
subject to $\bar{\boldsymbol{\Delta}}_{i}=\boldsymbol{\Delta}_{i}\quad\forall i$ (15)
$m_{i}\leq\delta_{t,i}\leq M_{i}\quad\forall t,i$ (16)